# The Kick Velocity Distribution of Isolated Neutron Stars

### v1

@@ Abstract
Neutron stars (NSs) are thought to receive natal kicks at their formation in supernovae. In order to investigate the magnitude of these kicks, we analyze the proper motions and distance estimates—either through parallax or dispersion measures—of young isolated pulsars and infer their three-dimensional velocities relative to their local standard of rest. We find that the velocities based on parallax distances of pulsars younger than Myr follow a log-normal distribution with a mean of and a standard deviation of , peaking at – km s-1, which we adopt as our fiducial kick distribution. Using a previously established method that infers kick magnitudes through the eccentricity of Galactic trajectories, we also estimate the kick velocities of older pulsars, which we find to be consistent with our fiducial kick distribution. A log-normal fit to all pulsars with ages below Myr yields a more constraining (but possibly more prone to systematic errors) fit with mean and standard deviation of and , respectively. Moreover, we (1) resolve the tension between our results and the Maxwellian distribution found by Hobbs et al. (2005), which has a higher median velocity, by showing that their analysis is missing a Jacobian needed to correct for its logarithmic histogram bin sizes, and (2) argue that the bimodality found by others is not statistically significant and that previous results are consistent with our inferred kick distribution, effectively reconciling the literature on observed NS kicks.
@@ 1 Introduction
Neutron stars (NSs) are born in supernovae which are likely anisotropic (e.g., Janka & Müller, 1994; Burrows et al., 1995), causing them to receive natal kick velocities (e.g., Van den Heuvel & Van Paradijs, 1997) with magnitudes of – km s-1 (Burrows et al., 2024). NSs observed as isolated pulsars are indeed known to have higher velocities than their progenitors (e.g., Gunn & Ostriker, 1970; Lyne & Lorimer, 1994), but the distribution of the kick velocities that these objects receive remains a topic of discussion. Several studies have tried to constrain the natal kicks of NSs based on the observed properties of isolated pulsars (e.g., Arzoumanian et al., 2002; Hobbs et al., 2005; Faucher-Giguère & Kaspi, 2006; Verbunt et al., 2017; Igoshev, 2020; Disberg et al., 2025), whereas other studies have investigated the natal kicks of NSs in binaries (e.g., Fortin et al., 2022; Zhao et al., 2023). These empirical natal kick distributions are often used in population synthesis models of binary evolution.
@@
For this reason, Verbunt et al. (2017) inferred the three-dimensional velocities of pulsars that have distance estimates determined through parallax (as opposed to DM), and found that these velocities are best described by a weighted sum of two Maxwellians. However, they note that this does not necessarily mean that the kick distribution is truly bimodal, since it could also be explained by a distribution that is wider than a single Maxwellian, and it is difficult to distinguish between these two cases since they fit their double-Maxwellian directly to the observational data instead of a posterior velocity distribution. Igoshev (2020), in turn, redid the analysis of Verbunt et al. (2017) after significantly expanding their pulsar sample with the parallax estimates of Deller et al. (2019), and also found a kick distribution best described by two Maxwellians (see also Igoshev et al., 2021), similarly to the results of Verbunt et al. (2017). The kick distributions of Verbunt et al. (2017) and Igoshev (2020) favor significantly lower velocities than the Maxwellian found by Hobbs et al. (2005).
@@ 2 Method
where and are the galactocentric cylindrical coordinates of the pulsar, where and for Galactic coordinates and galactocentric radius of the Solar System kpc (GRAVITY Collab. et al., 2018). We determine the parallax distances (for each parallax pulsar) by sampling values from Equation 1, whereas the DM distances (for all pulsars) are estimated by sampling values from a normal distribution centered around the distance determined through the electron density model of Yao et al. (2017), with a standard deviation of (Ding et al., 2024).
@@
where the transverse velocity equals and the radial velocity equals (note that “radial” here refers to the line of sight, rather than the Galactic radial coordinate). If this vector is oriented isotropically in the galactocentric (GC) frame, then is distributed uniformly and for . This means that for GC isotropy, the radial component of the velocity vector is given by
@@
We combine the distance estimates with the proper motions to create samples of the transverse velocity for each pulsar, and then use Equation 4 to create a corresponding samples for the radial velocity distribution, since LSR isotropy is a more accurate assumption than GC isotropy (Disberg et al., 2024b). The transverse and radial velocity samples form the three-dimensional velocity vectors in the galactocentric frame, which we transformed to velocities relative to the pulsars’ LSR by subtracting . We approximate by using the circular velocity at the location of the pulsar (as given by the Milky Way potential of McMillan, 2017). As a result we obtain the present-day velocities of the pulsars relative to their LSR—which we define as “”. For pulsars with characteristic spin-down ages below Myr, we approximate their kicks by estimating , since for pulsars with these ages their current velocities are still representative of their kicks (Disberg et al., 2024a). Our sample contains 142 pulsars with Myr, of which have parallax estimates. Moreover, although there are cases in which does not appear to be accurate (e.g., Lyne et al. 1996 find a factor – difference between and their age estimates), it tends to agree with kinematic age estimates (e.g., Noutsos et al., 2013; Igoshev, 2019; Disberg et al., 2025) and in particular Maoz & Nakar (2025) argue that it is a reasonable age estimate.
@@
The benefit of this method is that it can be used to obtain information about the kicks of old objects, whereas previous methods used to determine NS kicks only analyzed the current velocities of young objects. However, the downside of this method is that this model makes several assumptions that bring in more systematic uncertainties than the method used for young pulsars. For example, we integrate the eccentricity versus kick relationship over a time domain and compare it to pulsars with ages within this domain, but this relationship evolves over time and the distribution of pulsar ages is generally not uniform within a given time domain (see Appendix A). The boundaries of the time domains are sensitive to the accuracy of as an approximation for the pulsar age. Moreover, our model of the detection probability of a kicked neutron star, through the probability that it reaches (our definition of) the solar neighborhood, is an approximation that brings in systematic uncertainty as well. Because of these systematic model uncertainties, we base our fiducial kick distribution only on the young pulsars (i.e., Myr).
@@ 3 Results
In Table 1 we list the log-normal parameters (i.e., and ) fitted to pulsar velocities in the various age ranges. The younger the pulsars, the more representative their current velocities are for their natal kicks, but in order to limit the effects of low number statistics we consider the fit to pulsars with Myr (based on parallax distances), which results in and , to be our fiducial kick distribution. For the oldest pulsars with Myr, we find the data to be insufficient for constraining and . However, through Equation 6 we are able to fit a log-normal distribution to the pulsars with ages between and Myr, and by multiplying the likelihood for this sample with the likelihood for pulsars with Myr we constrain the parameters even further and obtain a fit for the pulsars with Myr, which results in and (for parallax distance estimates). Table 1 also lists fits based on DM distance estimates, which generally yield lower values of and higher values of , but we are more confident in the distance estimates based on parallax.
The statistical uncertainties on and quoted in Table 1 are half the width of the credible intervals on these quantities obtained from the posterior on and assuming the likelihood function of Equation 6 and flat priors. We note that for parallax distances our parameter estimates are consistent within their error bars, though the estimated for the sample is a marginal outlier. The parallax estimates are marginally consistent with the parameter estimates for DM distances, which favor smaller and wider .
@@ 4 Maxwellian
In other words, this histogram is missing the Jacobian and therefore shows instead of . We find that our histogram of that excludes this Jacobian is identical to the histogram of Hobbs et al. (2005, see their Figure 4b). Similarly, if we describe the distribution of three-dimensional velocities with a histogram that does not include a Jacobian and thus shows , the result is similar to the histogram of Hobbs et al. (2005, see their Figure 7).
However, to fit the histogrammed empirical distribution to the Maxwellian of Equation 7, a Jacobian must be included. We replot the histogram in uniform bins in panel B of Figure 2, including the Jacobian. It is apparent that the empirical distribution contains significantly lower velocities than predicted by the Hobbs et al. (2005) Maxwellian and is instead well-described by our log-normal fiducial distribution (despite the fact that these velocities are determined assuming GC isotropy instead of LSR isotropy).
@@ 5 Comparison
In panel D of Figure 3, we show CDFs of the present-day velocity distributions of these pulsar samples, for Myr. We estimate the confidence intervals by bootstrapping the distributions, similarly to our main results shown in Figure 1. We also show our fiducial kick distribution as well as our fit for Myr, both of which are encompassed by all three confidence intervals. We therefore argue that (1) correcting the results of Hobbs et al. (2005) for the logarithmic bin sizes reconciles them with the results of Verbunt et al. (2017) and Igoshev (2020), and (2) our fiducial kick distribution is compatible with all of these NS kicks analyses. Nevertheless, we do note that for larger sample sizes we expect the confidence regions of velocity distributions based on DM distances to be slightly below our kick distribution (as shown in Figures 1). Moreover, despite the fact that there are good reasons why the kick distribution may be bimodal (see Willcox et al., 2021; Disberg et al., 2025; Valli et al., 2025, and references therein), the bimodality found by Verbunt et al. (2017) and Igoshev (2020) is not statistically significant: the distributions are well described by our unimodal log-normal fiducial kick distribution. The bimodality is therefore likely a result of low-number statistics (e.g., the knee in the median CDF of the Verbunt et al. (2017) is no longer apparent in the bootstrapped 95% CDF confidence interval for the same sample).
@@ 6 Conclusions
The method we use to estimate the present-day pulsar velocities makes several assumptions. For instance, our fiducial kick distribution does not take into account a likely binary origin of the pulsars, in which the pulsars are kicked out of the binary system (e.g., Kuranov et al., 2009; Beniamini & Piran, 2024). This binary origin can impact the evolutionary history: e.g., binary interactions stripping supernova progenitors may affect kick magnitudes (Müller et al., 2018; Willcox et al., 2021). The pre-supernova orbital velocity of the binary will scatter the speed of the ejected single pulsar around the kick magnitude. Perhaps most importantly, binarity imposes a selection effect: only sufficiently large kicks will unbind binaries, potentially leading to an over-representation of high kicks among apparently single pulsars. However, Kapil et al. (2023) find that a binary origin does not significantly influence observed kicks in their population-synthesis models. We also do not consider the effects of overdensities in the Galactic disc on the velocities of kicked objects (Ofek, 2009) or the potential effect of a “rocket”-like acceleration due to spin-down radiation (Hirai et al., 2024). Lastly, we also do not consider a potential observational bias caused by spin-kick alignment (Biryukov & Beskin, 2024), which may cause pulsar velocities to be overestimated by (Mandel & Igoshev, 2023).
@@ Appendix D Maxwellian Fits
In Figure 3, we show the present-day velocity distributions for the pulsars in the samples of Igoshev (2020) and Verbunt et al. (2017), determined by assuming LSR isotropy (Equation 4) for all objects. The analyses of Igoshev (2020) and Verbunt et al. (2017) use a mixed model in which the radial velocities of a fraction of the sample are oriented semi-isotropically, based on the fact that pulsars are thought to be born in the thin disc of the Milky Way and young pulsars are therefore more likely to move away from the disc. Nevertheless, when we fit a double Maxwellian (Equation 9) to their data processed following our methodology, we find best-fit parameters within the uncertainties of their fits (with the slight exception of for the sample of Igoshev, 2020), as shown in Figure D. The main difference is that in our results the contribution of velocities above km s-1 is smaller, but this is likely not statistically significant. We therefore conclude that our method relatively accurately reproduces the results of Igoshev (2020) and Verbunt et al. (2017).

### v2

@@ Abstract
Neutron stars (NSs) are thought to receive natal kicks at their formation in supernovae. In order to investigate the magnitude of these kicks, we analyze the proper motions and distance estimates—either through parallax or dispersion measures—of young isolated pulsars and infer their three-dimensional velocities relative to their local standard of rest. We find that the velocities based on parallax distances of pulsars younger than Myr follow a log-normal distribution with and , peaking at – km s-1, which we adopt as our fiducial kick distribution. Using a previously established method that infers kick magnitudes through the eccentricity of Galactic trajectories, we also estimate the kick velocities of older pulsars, which we find to be consistent with our fiducial kick distribution. A log-normal fit to all pulsars with ages below Myr yields a more constraining (but possibly more prone to systematic errors) fit with and , respectively. Moreover, we (1) resolve the tension between our results and the Maxwellian distribution found by Hobbs et al. (2005), which has a higher median velocity, by showing that their analysis is missing a Jacobian needed to correct for its logarithmic histogram bin sizes, and (2) argue that the bimodality found by others is not statistically significant and that previous results are consistent with our inferred kick distribution, effectively reconciling the literature on observed NS kicks.
@@ 1 Introduction
Neutron stars (NSs) are born in supernovae which are likely anisotropic (e.g., Janka & Müller, 1994; Burrows et al., 1995; Herant, 1995), causing them to receive natal kick velocities (e.g., Van den Heuvel & Van Paradijs, 1997) with magnitudes of – km s-1 (Burrows et al., 2024). NSs observed as isolated pulsars are indeed known to have higher velocities than their progenitors (e.g., Gunn & Ostriker, 1970; Lyne & Lorimer, 1994), but the distribution of the kick velocities that these objects receive remains a topic of discussion. Several studies have tried to constrain the natal kicks of NSs based on the observed properties of isolated pulsars (e.g., Arzoumanian et al., 2002; Hobbs et al., 2005; Faucher-Giguère & Kaspi, 2006; Verbunt et al., 2017; Igoshev, 2020; Disberg et al., 2025), whereas other studies have investigated the natal kicks of NSs in binaries (e.g., Fortin et al., 2022; Zhao et al., 2023). These empirical natal kick distributions are often used in population synthesis models of binary evolution.
@@
For this reason, Verbunt et al. (2017) inferred the three-dimensional velocities of pulsars that have distance estimates determined through parallax (as opposed to DM), and found that these velocities are best described by a weighted sum of two Maxwellians. However, they note that this does not necessarily mean that the kick distribution is truly bimodal, since it could also be explained by a distribution that is wider than a single Maxwellian. Igoshev (2020), in turn, redid the analysis of Verbunt et al. (2017) after significantly expanding their pulsar sample with the parallax estimates of Deller et al. (2019), and also found a kick distribution best described by two Maxwellians (see also Igoshev et al., 2021), similarly to the results of Verbunt et al. (2017). The kick distributions of Verbunt et al. (2017) and Igoshev (2020) favor significantly lower velocities than the Maxwellian found by Hobbs et al. (2005).
@@ 2 Method
where and are the galactocentric cylindrical coordinates of the pulsar, where and for Galactic coordinates and galactocentric radius of the Solar System kpc (GRAVITY Collab. et al., 2018). We determine the parallax distances (for each parallax pulsar) by sampling values from Equation 1, whereas the DM distances (for all pulsars) are estimated by sampling values from a normal distribution centered around the distance determined through the electron density model of Yao et al. (2017), with a standard deviation of . Although detailed studies such as those of Deller et al. (2009) show that DM distances can sometimes have even greater systematic errors, we use an indicative error of (e.g. Ding et al., 2024). We note that we primarily use DM distance estimates to show that these are not sufficient to explain the differences between the results of Verbunt et al. (2017) and those of Hobbs et al. (2005).
@@
where the transverse velocity equals and the radial velocity equals (note that “radial” here refers to the line of sight, rather than the Galactic radial coordinate). If this vector is oriented isotropically in the galactocentric frame (GC isotropy), then is distributed uniformly and for . This means that for GC isotropy, the radial component of the velocity vector is given by
@@
We combine the distance estimates with the proper motions to create samples of the transverse velocity for each pulsar, and then use Equation 4 to create a corresponding sample for the radial velocity distribution, since LSR isotropy is a more accurate assumption than GC isotropy (Disberg et al., 2024b). The transverse and radial velocity samples form the three-dimensional velocity vectors in the galactocentric frame, which we transformed to velocities relative to the pulsars’ LSR by subtracting . We approximate by using the circular velocity at the location of the pulsar (as given by the Milky Way potential of McMillan, 2017). As a result we obtain the present-day velocities of the pulsars relative to their LSR—which we define as “”. For pulsars with characteristic spin-down ages below Myr, we approximate their kicks by estimating , since for pulsars with these ages their current velocities are still representative of their kicks (Disberg et al., 2024a). Our sample contains 142 pulsars with Myr, of which have parallax estimates. Moreover, although there are cases in which does not appear to be accurate (e.g., Lyne et al. 1996 find a factor – difference between and their age estimates), it tends to agree with kinematic age estimates (e.g., Noutsos et al., 2013; Igoshev, 2019; Disberg et al., 2025) and in particular Maoz & Nakar (2025) argue that it is a reasonable age estimate.
@@
The benefit of this method is that it can be used to obtain information about the kicks of old objects, whereas previous methods used to determine NS kicks only analyzed the current velocities of young objects. However, the downside of this method is that this model makes several assumptions that bring in more systematic uncertainties than the method used for young pulsars. For example, we integrate the eccentricity versus kick relationship over a time domain and compare it to pulsars with ages within this domain, but this relationship evolves over time and the distribution of pulsar ages is generally not uniform within a given time domain (see Appendix A). The boundaries of the time domains are sensitive to the accuracy of as an approximation for the pulsar age. We assume a static Milky Way model because Galactic evolution since the formation of the pulsars is not significant (e.g. Amend et al., 2024). Moreover, our model of the detection probability of a kicked neutron star, through the probability that it reaches (our definition of) the solar neighborhood, is an approximation that brings in systematic uncertainty as well. Because of these systematic model uncertainties, we base our fiducial kick distribution only on the young pulsars (i.e., Myr).
@@ 3 Results
In Table 1 we list the log-normal parameters (i.e., and ) fitted to pulsar velocities in the various age ranges. The younger the pulsars, the more representative their current velocities are for their natal kicks, but in order to limit the effects of low number statistics we consider the fit to pulsars with Myr (based on parallax distances), which results in and , to be our fiducial kick distribution. For the oldest pulsars with Myr, we find the data to be insufficient for constraining and . However, through Equation 6 we are able to fit a log-normal distribution to the pulsars with ages between and Myr, and by multiplying the likelihood for this sample with the likelihood for pulsars with Myr we constrain the parameters even further and obtain a fit for the pulsars with Myr, which results in and (for parallax distance estimates).
The statistical uncertainties on and quoted in Table 1 are half the width of the credible intervals on these quantities obtained from the posterior on and assuming the likelihood function of Equation 6 and flat priors. We note that for parallax distances our parameter estimates are consistent within their error bars, though the estimated for the sample is a marginal outlier. Table 1 also lists fits based on DM distance estimates, but we are more confident in the distance estimates based on parallax. The parallax estimates are marginally consistent with the parameter estimates for DM distances, which favor smaller and wider .
@@
In our fiducial kick distribution only of pulsars obtain low kicks (below km s-1, following Willcox et al., 2021), which are relevant for retaining NSs in globular clusters or wide binaries, although in the confidence region this fraction may be as large as . It seems likely that there is a subpopulation of NSs that are formed following binary interactions, receive low natal kicks, therefore remain in binaries such as low-eccentricity Be X-ray binaries (e.g., Valli et al., 2025), wide low-eccentricity detached NS binaries observed with Gaia (though post-natal kick mechanisms could be at play, see Hirai et al. 2024) and in clusters (e.g., Smits et al., 2006), and are thus explicitly excluded from our observational isolated pulsar data set. These might, for example, originate from electron-capture supernovae that happen predominantly or exclusively for progenitors stripped by mass transfer and thus avoiding second dredge-up (Willcox et al., 2021) or from accretion induced collapse. Another possibility to consider is that NSs that receive low kicks are somehow different from typical pulsars and avoid pulsar surveys unless they are recycled, but there is no evidence for a correlation between kick velocity and pulsar detectability in the existing data (Willcox et al., 2021).
@@ 4 Maxwellian
In other words, this histogram is missing the Jacobian and therefore shows instead of . We find that our histogram of that excludes this Jacobian is identical to the histogram of Hobbs et al. (2005, see their Figure 4b).
Similarly, if we describe the distribution of three-dimensional velocities with a histogram that does not include a Jacobian and thus shows , the result is similar to the histogram of Hobbs et al. (2005, see their Figure 7). However, to fit the histogrammed empirical distribution to the Maxwellian of Equation 7, a Jacobian must be included. We replot the histogram in uniform bins in panel B of Figure 2, including the Jacobian. It is apparent that the empirical distribution contains significantly lower velocities than predicted by the Hobbs et al. (2005) Maxwellian and is instead well-described by our log-normal fiducial distribution (despite the fact that these velocities are determined assuming GC isotropy instead of LSR isotropy).
@@ 5 Comparison
In panel D of Figure 3, we show CDFs of the present-day velocity distributions of these pulsar samples, for Myr. We estimate the confidence intervals by bootstrapping the distributions, similarly to our main results shown in Figure 1. We also show our fiducial kick distribution as well as our fit for Myr, both of which are encompassed by all three confidence intervals. We therefore argue that (1) correcting the results of Hobbs et al. (2005) for the logarithmic bin sizes reconciles them with the results of Verbunt et al. (2017) and Igoshev (2020), and (2) our fiducial kick distribution is compatible with all of these NS kicks analyses. Nevertheless, we do note that for larger sample sizes we expect the confidence regions of velocity distributions based on DM distances to be slightly below our kick distribution (as shown in Figures 1).
Moreover, despite the fact that there are good reasons why the kick distribution may be bimodal (see e.g. Katz, 1975; Fryer et al., 1998; Podsiadlowski et al., 2004; Fryer & Young, 2007; Banerjee et al., 2020; Willcox et al., 2021; Disberg et al., 2025; Valli et al., 2025, and references therein), the bimodality found by Verbunt et al. (2017) and Igoshev (2020) is not statistically significant: the distributions are well described by our unimodal log-normal fiducial kick distribution. The bimodality is therefore likely a result of low-number statistics. In order to examine this, we randomly select pulsars (the number of young pulsars used in the Igoshev 2020 analysis) from our sample of Myr parallax pulsars and evaluate the histogrammed velocity PDFs. We find that the PDF drops below s km-1 in the 200–250 km s-1 in 10% of the 1000 random draws of pulsars, showing a bimodality similar to the distributions of Verbunt et al. (2017) and Igoshev (2020). Lastly, we investigate whether a distribution such as a Maxwellian or a (bimodal) double-Maxwellian could provide a better fit to the data than our log-normal model. As shown in Appendix E, we do not find strong evidence in favor of an alternative kick distribution.
@@ 6 Conclusions
The method we use to estimate the present-day pulsar velocities makes several assumptions. For instance, our fiducial kick distribution does not take into account a likely binary origin of the pulsars, in which the pulsars are kicked out of the binary system (e.g., Kuranov et al., 2009; Beniamini & Piran, 2024). This binary origin can impact the evolutionary history: for example, binary interactions stripping supernova progenitors may affect kick magnitudes (Müller et al., 2018; Willcox et al., 2021). Also, the pre-supernova orbital velocity of the binary will scatter the speed of the ejected single pulsar around the kick magnitude. Perhaps most importantly, binarity imposes a selection effect: only sufficiently large kicks will unbind binaries, potentially leading to an over-representation of high kicks among apparently single pulsars. However, Kapil et al. (2023) find that a binary origin does not significantly influence observed kicks in their population-synthesis models. Moreover, we neglect the effects of overdensities in the Galactic disc on the velocities of kicked objects (Ofek, 2009) or the potential effect of a “rocket”-like acceleration due to spin-down radiation (Hirai et al., 2024). Lastly, we also do not consider a potential observational bias caused by spin-kick alignment (Biryukov & Beskin, 2024), which may cause pulsar velocities to be overestimated by (Mandel & Igoshev, 2023).
@@ Appendix D Maxwellian Fits
In Figure 3, we show the present-day velocity distributions for the pulsars in the samples of Igoshev (2020) and Verbunt et al. (2017), determined by assuming LSR isotropy (Equation 4) for all objects. The analyses of Igoshev (2020) and Verbunt et al. (2017) use a mixed model in which the radial velocities of a fraction of the sample are oriented semi-isotropically, based on the fact that pulsars are thought to be born in the thin disc of the Milky Way and young pulsars are therefore more likely to move away from the disc. Nevertheless, when we fit a double-Maxwellian (Equation 9) to their data processed following our methodology, we find best-fit parameters within the uncertainties of their fits (with the slight exception of for the sample of Igoshev, 2020), as shown in Figure D. The main difference is that in our results the contribution of velocities above km s-1 is smaller, but this is likely not statistically significant. We therefore conclude that our method relatively accurately reproduces the results of Igoshev (2020) and Verbunt et al. (2017).
Appendix E Model Comparison
Our fiducial kick distribution is a log-normal distribution fitted to the pulsar velocity estimates from our sample, and in Figure 1 we show that this fit stays within the bootstrapped confidence region of the observed CDFs. However, the question arises whether an alternative model, such as a double-Maxwellian model, might be a better description of our results. We therefore fit several distributions to our results for the parallax pulsars with Myr: a Maxwellian, a double-Maxwellian, and a Weibull distribution—a unimodal distribution defined as:
where the parameters and determine the shape and scale of the distribution, respectively.
For these fitted distributions, as well as our log-normal fiducial kick distribution and the literature distributions of Hobbs et al. (2005), Verbunt et al. (2017) and Igoshev (2020), we evaluate the Bayesian information criterion (BIC), defined as:
where equals the number of parameters in the distribution function, is the number of pulsars in the sample, and is the maximum likelihood (Equation 6). Kass & Raftery (1995) state that a difference in BIC between and corresponds to “strong” evidence in favor of the model with the lowest BIC, and if the difference exceeds the evidence is “decisive”.
In Table E, we list the fitted parameters of the considered distributions, as well as the values of the BIC and the difference between their BIC and the BIC of our fiducial kick distribution (BIC). These distributions are shown in Figure E. For the fitted Maxwellian, double-Maxwellian, and Weibull distribution, the value of is less than , meaning there is no strong evidence in favor of either distribution. We find that hypothetical pulsars with no measurement uncertainty with kicks sampled from the Igoshev (2020) distribution would be sufficient to obtain between the fitted log-normal and double-Maxwellian distributions, meaning that the small BIC values are mainly caused by observational uncertainties and the necessity to estimate (Equation 4). Although the Maxwellian has a significantly lower likelihood, it has the benefit of having only one parameter. Moreover, the best-fit double-Maxwellian to our larger sample is significantly less bimodal than the distributions shown in Figure D. We note that if we fit a scaled beta distribution (O’Doherty et al., 2023) or an asymmetric Gaussian (Disberg & Nelemans, 2023), the resulting BIC values are also within the range where these functional forms are neither preferred nor ruled out given the available data. Lastly, the literature distributions, taking the best-fit values from the literature,333We still account for the full number of parameters for the relevant distributions when computing BICs given the largely overlapping data sets, though one could arguably penalize these distributions less given that there are no free parameters in the quoted fits. have higher BIC values, where there is strong evidence against the (erroneous) distribution of Hobbs et al. (2005) and decisive evidence against the distribution of Igoshev (2020).
| distribution | fit | BIC | BIC |
| log-normal | 666.2 | 0 | |
| Maxwellian | 666.4 | 0.2 | |
| double-Maxwellian | 665.7 | -0.5 | |
| Weibull distribution | 662.7 | -3.5 | |
| Hobbs et al. (2005) | 673.1 | 6.9 | |
| Verbunt et al. (2017) | 671.3 | 5.1 | |
| Igoshev (2020) | 678.2 | 12 |
The values in parentheses are half the widths of the credible intervals. The log-normal distribution is our fiducial kick distribution (Table 1). The value of BIC equals the BIC of the distribution minus the BIC of our fiducial kick distribution. The parameters , , , and are given in km s-1.